OpenStudy (anonymous):

HOW SHOULD I DO THIS PROBLEM????? x + y = k 2x + 3y = k + 1 The point of intersection of the lines has an x-coordinate of 2k + 1 -2k + 1 2k - 1

OpenStudy (anonymous):

multiply the top equation by -3 to eliminate the y's giving you -3x - 3y = -3k 2x + 3y = k + 1 -x + 0y = -2k + 1 -x = -2k + 1 x = 2k - 1

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

hey, how did u come up with -3????

OpenStudy (anonymous):

because you are only wanting the x coordinate so you have to make y go away. And the only way to do that is to multiply the top equation by -3 so when you add the two equations together your x's are still there but you get a 0*y and zero times anything is 0 so the y's cancel which is what you want.

OpenStudy (anonymous):

ok thankx for the help

OpenStudy (anonymous):

how about this one???? Solve the system by the elimination method. 2x + y - 4 = 0 2x - y - 4 = 0 When you eliminate y, what is the resulting equation? x = 0 4x = -8 4x = 8